A cannon of mass M starts sliding freely down a smooth inclined plane at an angle α to the horizontal. After the cannon covered the distance l, a shot was fired, the shell leaving the cannon in the horizontal direction with a momentum p. As a consequence, the cannon stopped. Assuming the mass of the shell to be negligible, as compared to that of the cannon, determine the duration of the shot.
1
Expert's answer
2020-10-15T10:43:56-0400

Solution

According to question velocity of cannon is"\sqrt{2gl \sin\alpha}" down to inclined plane. using linear impulse momentum theorem

"\Delta p_x=F_xdt"

"\implies p\cos \alpha-M\sqrt{2gl\sin\alpha}"

"=Mg\sin \alpha\Delta t"


"\Delta t=\frac{p\cos \alpha-M\sqrt{2gl\sin\alpha}}{Mg\sin \alpha}"